OpenStudy (anonymous):

-2x+4=2(4x-3)-3(-8+4x)

OpenStudy (john_es):

Simplify, using the distributive property, \[−2x+4=8x−6+24−12x\] Group numbers in one side and variables in other, \[4+6−24=8x+2x−12x⇒−14=−2x⇒x=7\] Sorry typing problems.

OpenStudy (anonymous):

See more at: http://www.tiger-algebra.com/drill you will get your answer here ^.^

OpenStudy (john_es):

Good link :)

OpenStudy (anonymous):

yw lol

OpenStudy (anonymous):

oh if you type your problem into google it will come up with answers lol im in internet school so google is my text book

OpenStudy (anonymous):

know of something that would help with this problem? match the inequality to the graph of its solution